# Enhancing Financial Time-Series Forecasting with Retrieval-Augmented Large Language Models

### v2

@@ 4.3. Ablation Study
As shown in the figure, FinSeer is the only model that successfully extracts a diverse and comprehensive set of indicators while achieving superior performance. This clearly demonstrates its advanced temporal retrieval capabilities. Specifically, while other models like LLM Embedder and Instructor predominantly focus on basic indicators such as close price and adjusted close price, FinSeer effectively identifies and retrieves a wide range of technical indicators, including kdj crossover, MACD Histogram, Bollinger Bands, and various alpha factors. This richer set of retrieved indicators provides FinSeer with more comprehensive auxiliary information, enabling more accurate and reliable predictions.
@@
Candidate Movement Correlation In this section, we investigate whether the LLM relies on the movement trends of retrieved sequences to make predictions. To this end, we compute the correlation between the movements of retrieved sequences and the LLM’s generated results. The detailed calculation process is provided in Appendix B.2.
@@
The results in Table 4 reveal no significant correlation between the movement direction (rise or fall) of the retrieved candidate sequences and the final predictions of StockLLM. This finding indicates that StockLLM does not simply mirror the movement trends of the retrieved sequences but instead analyzes their specific content to infer the query’s movement direction. This ability highlights the LLM’s capacity to extract meaningful insights from complex time-series data.
@@
Table 4. Correlation between the movements of retrieved sequences and LLM’s generated results.
@@
This observation underscores the critical role of retrieval quality in RAG models, as the relevance and informativeness of retrieved sequences directly influence prediction performance. Notably, StockLLM is a relatively small 1B parameter LLM, making this a particularly challenging setting. Despite this, our time-series-tailored RAG model, FinSeer, successfully enhances StockLLM’s performance by retrieving rich and relevant time-sensitive information. This demonstrates the effectiveness of our approach in improving prediction accuracy and showcases its potential for financial time-series forecasting.
4.4. Case Study
{ "query_stock": "XOM",
"query_date": "2015-06-25",
While multiple retrievers were evaluated, only FinSeer successfully enabled StockLLM to predict the correct movement as a fall. Specifically, FinSeer retrieved five diverse indicators—close price, adjusted close price, alpha021, alpha054, and the highest price—providing a comprehensive view of the stock’s behavior. Alpha021 identifies trends based on short- and long-term price averages and volume conditions, while alpha054 combines price and volume rankings to assess performance within a specific time window. These indicators allowed StockLLM to accurately assess whether the downward trend would persist or reverse, demonstrating the value of retrieving contextually relevant and diverse features.
4.5. Indicator Sequence Visualization
@@
To intuitively examine how FinSeer embeds indicator sequences, we use LargeVis (Tang et al., 2016) to reduce the embeddings to two dimensions. In Figure 3, we visualize [^3] https://medviz.org/app/ the vector space of nine indicators, including adjusted close price, open price, high price, low price, close price, volume, movement, returns, and WAP. The results show that the green points, representing movement information, are well-clustered, indicating that FinSeer effectively captures meaningful patterns. However, this also suggests that deeper relationships between stock movements and indicators remain to be explored, highlighting the potential for further research in this area.
Figure 3. Visualization of indicator sequence embeddings on ACL18 dataset.
[…]
5. Related Work
5.1. Stock Movement Prediction
Non-LLM Methods. Traditional approaches to stock movement prediction have focused on various aspects of financial data. One prominent category of methods analyzes stock price sequences and their corresponding technical indicators (Qin et al., 2017; Feng et al., 2018) to identify patterns in historical data for predicting future movements. However, due to the complexity of factors influencing stock prices, subsequent methods have incorporated additional contextual information, such as news articles (Ding et al., 2015; Liu et al., 2018) or social media posts (Xu and Cohen, 2018; Wu et al., 2018). Despite efforts, these methods are highly susceptible to noise and struggle to analyze the vast and diverse nature of financial information effectively.
LLM-based Methods. Recent studies have explored using LLMs for financial prediction tasks, either by fine-tuning open-source models or prompting advanced models like GPT-4. However, even state-of-the-art models, including GPT-4, have achieved only random-guessing-level accuracy in stock movement prediction (Xie et al., 2023b, 2024a, a). This highlights the inherent challenges in identifying and analyzing meaningful patterns in a domain as volatile and multifaceted as stock market prediction.
5.2. Time-series forecasting with LLMs
Non-retrieval Methods. To enhance the performance of LLMs in time-series forecasting, existing methods have primarily focused on aligning temporal and textual data, either by transforming time-series into textual formats or by encoding both modalities into a unified vector space. For instance, a study (Jin et al., 2023) reprograms time-series data into textual representations suitable for LLMs, enhancing prediction accuracy via declarative prompts. Similarly, some other studies (Yu et al., 2023; Liu et al., 2024) explore cross-modal alignment: the former applies LLMs to financial forecasting by integrating stock prices with news data, while the latter introduces a cross-modality framework to align time-series with text for improved predictive performance. While these advancements improve alignment, existing methods often struggle to process large-scale time-series data comprehensively due to LLM input limitations. As a result, effectively leveraging extensive historical data remains a challenge. This highlights the need for retrieval-augmented methods, a gap our approach directly addresses.
RAG Methods. With the rapid advancement of RAG techniques in various applications of LLMs, recent research has begun exploring RAG for time-series forecasting. For instance, TimeRAG (Yang et al., 2024) integrates RAG into time-series forecasting by combining Dynamic Time Warping (DTW) with LLMs to improve prediction accuracy. However, relying solely on numeric similarity is insufficient for financial time-series forecasting, as it fails to capture deeper semantic relationships. This underscores the need for a more targeted retrieval framework tailored to the complexities of financial data, a challenge our framework effectively addresses.
6. Conclusion
In this work, we introduce the first RAG framework for financial time-series forecasting, featuring a novel retrieval mechanism with three key innovations: a fine-tuned 1B parameter LLM (StockLLM), a candidate selection method leveraging LLM feedback, and a training objective that maximizes similarity between queries and historically significant sequences. We also construct new datasets integrating financial indicators and historical stock prices, selected through expert consultation and mutual information analysis, to train the retriever and ensure robust evaluation. Our retriever, FinSeer, leverages these datasets and LLM feedback to uncover deeper patterns beyond numeric similarities. Experiments demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, while FinSeer surpasses existing retrieval methods, achieving an 8% higher accuracy on BIGDATA22 and retrieving more impactful sequences. This work highlights the importance of tailored retrieval models in financial forecasting and provides a robust framework for future research.
Appendix A Data examples
A.1. Examples in the Datastore
An example in our datastore is shown below.
"stock_name":"ABBV",
"query_date":"2014-06-05",
"query_movement":"rise",
"open":54.549999,
"high":55.32,
"low":54.360001,
"close":55.299999,
"adj_close":36.980961,
"volume":4847300,
"MACD_Histogram":0.0621141602,
"macd_crossover":null,
"bollinger_bands":null,
"exceeding_upper":null,
"exceeding_lower":null,
"overbought_and_oversold_conditions":null,
"kdj_crossover":null,
"Returns":0.0131913094,
"VWAP":54.9933333333,
"alpha_smr":0.000549858,
"alpha_mom":0.0121661224,
"alpha002":null,
"alpha006":null,
"alpha009":-0.481476,
"alpha012":-0.481476,
"alpha021":1,
"alpha023":0.0,
"alpha024":-0.769035,
"alpha028":null,
"alpha032":null,
"alpha041":-0.1554335256,
"alpha046":-0.481476,
"alpha049":-0.481476,
"alpha051":-0.481476,
"alpha053":null,
"alpha054":126.4264005641,
"alpha101":-18.282056485
Appendix B Addtional Experimental Results
B.1. Score Calculation of Financial Indicators
We compute the MI scores for each indicator using nonparametric estimation methods provided by mutual_info_regression from the Scikit-learn library. Then we normalize them into range $(0,1)$. These scores provide insight into the strength of the dependency between each financial indicator and the forward return. Higher MI scores indicate stronger relationships, highlighting which indicators are most predictive of future price movements. We select the top-20 indicator groups with the highest MI scores in our candidates, as is shown in Figure 4.
Figure 4. Mutual Information(MI) scores between indicators and forward return.
B.2. Candidate Movement Correlation
For each query with retrieved candidates, we convert the movements of retrieved sequences and LLM’s prediction to numerical representations.
[…]

### v3

@@ 4.3 Ablation Study
In this section, we analyze the retrieved indicators of all RAG models. We calculate indicator occurrences on the ACL18 test set, and the results are shown in Figure 3. As shown in the figure, FinSeer is the only model that successfully extracts a diverse and comprehensive set of indicators while achieving superior performance. This clearly demonstrates its advanced temporal retrieval capabilities. Specifically, while other models like LLM Embedder and Instructor predominantly focus on basic indicators such as close price and adjusted close price, FinSeer effectively identifies and retrieves a wide range of technical indicators, including kdj crossover, MACD Histogram, Bollinger Bands, and various alpha factors. This richer set of retrieved indicators provides FinSeer with more comprehensive auxiliary information, enabling more accurate and reliable predictions.
Figure 3: Indicator occurrences of different RAG models on ACL18 dataset.
4.4 Case Study
{ "query_stock": "XOM",
"query_date": "2015-06-25",
While multiple retrievers were evaluated, only FinSeer successfully enabled StockLLM to predict the correct movement as a fall. Specifically, FinSeer retrieved five diverse indicators: close price, adjusted close price, alpha021, alpha054, and the highest price, providing a comprehensive view of the stock’s behavior. Alpha021 identifies trends based on short- and long-term price averages and volume conditions, while alpha054 combines price and volume rankings to assess performance within a specific time window. These indicators allowed StockLLM to accurately assess whether the downward trend would persist or reverse, demonstrating the value of retrieving contextually relevant and diverse features.
5 Conclusion
In this paper, we present FinSrag, the first retrieval-augmented generation framework tailored for financial time-series forecasting. At its core, FinSeer, a novel retriever refined by LLM feedback, effectively identifies historically influential market sequences while filtering out financial noise. Combined with StockLLM, a fine-tuned LLM with 1B parameters, our framework leverages enriched financial datasets to capture previously overlooked market dynamics. Empirical results confirm that FinSeer surpasses both textual and distance-based retrievers in improving StockLLM’s prediction accuracy, highlighting the necessity of domain-specific retrieval in financial forecasting. Beyond stock markets, FinSrag establishes a blueprint for integrating RAG in time-sensitive decision-making domains.
Appendix A Limitation
While our FinSrag framework demonstrates significant improvements in LLM-based time-series forecasting, several limitations suggest important directions for future research. First, although our current evaluation focuses specifically on stock movement prediction, the core methodology and architectural principles of our framework are inherently generalizable to other time-series prediction tasks. Future work will systematically validate this generalizability across domains, including but not limited to economic indicators, energy demand forecasting, and epidemiological trend prediction.
Second, the present study deliberately focuses on unimodal time-series data to provide clear, controlled evidence for the effectiveness of our retrieval mechanisms. This focused approach has allowed us to rigorously quantify the contribution of temporal pattern retrieval independent of other factors. However, we recognize that many real-world forecasting scenarios could benefit from multimodal integration, where time-series data interacts with textual reports, visual data, or other information modalities. Future work will incorporate cross-modal corpora to enable more comprehensive information retrieval.
Appendix B Ethical Statement
Our framework and its contents are provided for academic and educational purposes only. None of the material constitutes financial, legal, or investment advice. No warranties, express or implied, are offered regarding the accuracy, completeness, or utility of the content. The authors and contributors are not responsible for any errors, omissions, or any consequences arising from the use of the information herein. Users should exercise their own judgment and consult professionals before making any financial, legal, or investment decisions. The use of the software and information contained in this repository is entirely at the user’s own risk.
FinSrag and all its components are licensed under MIT.
Appendix C Task Definition Details
[…]
Following Yoo et al. [2021] and  Soun et al. [2022], we classify the movement as rise if return $R_{t}$ exceeds 0.55, fall if it is below -0.5. In line with previous experimental settings [Xie et al., 2023b, 2024b], we do not evaluate freeze cases as queries. However, we include sequences with $R_{t}\in[-0.5,0.55]$ as freeze candidates in the candidate pool, ensuring a diverse and comprehensive set of historical data for context.
Appendix D Related Work
D.1 Stock Movement Prediction
Non-LLM Methods. Traditional approaches to stock movement prediction have focused on various aspects of financial data. One prominent category of methods analyzes stock price sequences and their corresponding technical indicators [Qin et al., 2017, Feng et al., 2018] to identify patterns in historical data for predicting future movements. However, due to the complexity of factors influencing stock prices, subsequent methods have incorporated additional contextual information, such as news articles [Ding et al., 2015, Liu et al., 2018] or social media posts [Xu and Cohen, 2018, Wu et al., 2018]. Despite efforts, these methods are highly susceptible to noise and struggle to analyze the vast and diverse nature of financial information effectively.
LLM-based Methods. Recent studies have explored using LLMs for financial prediction tasks, either by fine-tuning open-source models or prompting advanced models like GPT-4. However, even state-of-the-art models, including GPT-4, have achieved only random-guessing-level accuracy in stock movement prediction [Xie et al., 2023b, 2024b, a]. This highlights the inherent challenges in identifying and analyzing meaningful patterns in a domain as volatile and multifaceted as stock market prediction.
D.2 Time-series forecasting with LLMs
Non-retrieval Methods. To enhance the performance of LLMs in time-series forecasting, existing methods have primarily focused on aligning temporal and textual data, either by transforming time-series into textual formats or by encoding both modalities into a unified vector space. For instance, a study [Jin et al., 2023] reprograms time-series data into textual representations suitable for LLMs, enhancing prediction accuracy via declarative prompts. Similarly, some other studies [Yu et al., 2023, Liu et al., 2024] explore cross-modal alignment: the former applies LLMs to financial forecasting by integrating stock prices with news data, while the latter introduces a cross-modality framework to align time-series with text for improved predictive performance. While these advancements improve alignment, existing methods often struggle to process large-scale time-series data comprehensively due to LLM input limitations. As a result, effectively leveraging extensive historical data remains a challenge. This highlights the need for retrieval-augmented methods, a gap our approach directly addresses.
RAG Methods. With the rapid advancement of RAG techniques in various applications of LLMs, recent research has begun exploring RAG for time-series forecasting. For instance, TimeRAG [Yang et al., 2024] integrates RAG into time-series forecasting by combining Dynamic Time Warping (DTW) with LLMs to improve prediction accuracy. However, relying solely on numeric similarity is insufficient for financial time-series forecasting, as it fails to capture deeper semantic relationships. This underscores the need for a more targeted retrieval framework tailored to the complexities of financial data, a challenge our framework effectively addresses.
Appendix E Dataset Construction Details
E.1 Dataset Partition
To construct our datasets, we select high-trade-volume U.S. stocks across three periods: 2014-2015, 2017-2018, and 2022-2023. The first two periods align with two benchmark datasets: for 2014-2015, we use the same stocks as the ACL18 dataset [Xu and Cohen, 2018], and for 2017-2018, we use the same stocks as the BIGDATA22 dataset [Soun et al., 2022]. To incorporate more recent data, we manually retrieve high-trade-volume stocks from 2022 to 2023 to create the STOCK23 dataset. This ensures our training and evaluation reflect recent market conditions and provide a robust benchmark for stock movement prediction.
To maintain temporal integrity and prevent information leakage, we implement stock-wise partitioning rather than temporal splitting of the dataset. This approach addresses a critical characteristic of our data where many high-trade-volume stocks (e.g., Google (GOOG)) appear across multiple years (2014-2015 and 2019-2020).
Table 3 details the complete stock lists and their corresponding sectors for all subsets, demonstrating two key properties of our partitioning: first, the complete absence of overlap between training and test sets, and second, comprehensive sector coverage in both sets. The alphabetical partitioning serves as a neutral, systematic approach that is uncorrelated with any financial or temporal characteristics while achieving our core objectives of clean separation and representative sampling. This method guarantees that the model encounters truly novel stocks during testing without any temporal leakage, while maintaining balanced sector representation across all partitions.
Table 3: Stock information in datasets.
Dataset Split (amount) Stock List Sector Count
train (33) ABBV, AEP, AMGN, AMZN, BA, BABA, BAC, BCH, BHP, BP, BRK-A, BSAC, BUD, C, CAT, CHTR, CMCSA, CODI, CSCO, CVX, D, DHR, DIS, DUK, EXC, GD, GE, GOOG, HD, HON, HSBC, IEP, INTC basic-materials (1), communication-services (4), consumer-cyclical (3), consumer-defensive (1), energy (3), financial-services (6), healthcare (3), industrials (6), technology (2), utilities (4)
ACL18 valid (5) JNJ, JPM, KO, LMT, MA consumer-defensive (1), financial-services (2), healthcare (1), industrials (1)
test (33) MCD, MDT, MMM, MO, MRK, MSFT, NEE, NGG, NVS, ORCL, PCG, PEP, PFE, PG, PM, PPL, REX, SLB, SNY, SO, SPLP, SRE, T, TM, TSM, UL, UNH, UPS, V, VZ, WFC, WMT, XOM basic-materials (1), communication-services (2), consumer-cyclical (2), consumer-defensive (6), energy (2), financial-services (2), healthcare (6), industrials (3), technology (3), utilities (6)
train (22) AEP, AMGN, AMZN, BA, BAC, C, CAT, CODI, CSCO, CVX, D, DIS, DUK, EXC, GD, GE, GMRE, GOOG, HD, HON, INTC, JNJ communication-services (2), consumer-cyclical (2), energy (1), financial-services (2), healthcare (2) industrials (6), real-estate (1), technology (2), utilities (4)
BIGDATA22 valid (3) JPM, KO, LMT financial-services (1), consumer-defensive (1), industrials (1)
test (22) MA, MCD, MDT, MMM, MO, MRK, MSFT, NEE, ORCL, PCG, PM, PPL, REX, SO, SRE, T, UPS, V, VZ, WFC, WMT, XOM basic-materials (1), communication-services (2), consumer-cyclical (1), consumer-defensive (3), energy (1), financial-services (3), healthcare (2), industrials (2), technology (2), utilities (6)
train (24) ADBE, ADSK, AMD, AVGO, COIN, COST, DELL, DXCM, ELV, ETSY, FOX, HOOD, HPQ, IBM, INTU, JBL, JNJ, KKR, KMX, LDOS, LLY, LOW, LULU, LUV communication-services (1), consumer-cyclical (4), consumer-defensive (1), financial-services (3), healthcare (4), industrials (1), technology (10)
STOCK23 valid (3) MARA, META, MNST financial-services (1), communication-services (1), consumer-defensive (1)
test (24) MRVL, MS, MSCI, NDAQ, NKE, NVDA, ORLY, PANW, PEAK, PLTR, PNC, PNW, RY, SMCI, SNOW, SNPS, T, TGT, TSLA, UBER, ULTA, V, WBA, WDAY communication-services (1), consumer-cyclical (4), consumer-defensive (1), financial-services (6), healthcare (2), technology (9), utilities (1)
E.2 Financial Indicators
Financial Indicators Selection We compute the MI scores for each indicator using nonparametric estimation methods provided by mutual_info_regression from the Scikit-learn library. For each query with retrieved candidates, we convert the movements of retrieved sequences and LLM’s prediction to numerical representations.
[…]
Then we normalize them into range $(0,1)$. These scores provide insight into the strength of the dependency between each financial indicator and the forward return. Higher MI scores indicate stronger relationships, highlighting which indicators are most predictive of future price movements. We select the top-18 alpha indicators with the highest MI scores in our candidates, as is shown in Figure 4.
Figure 4: Mutual Information(MI) scores between indicators and forward return.
Indicators Descriptions We provided a brief description of each indicator in Table 4.
Table 4: Brief descriptions of indicators.
Group Indicator Description
basic price indicators
1 open (opening price) CoinAPI [2023] reflecting the stock’s value at the start of the trading session.
high, low (highest/lowest price) CoinAPI [2023] capture the price range during the day, offering insights into volatility.
close (close price) CoinAPI [2023] often the most important for technical analysis, represents the final trading price for the day.
adj_close (adjusted close price) Ganti [2020] accounts for corporate actions like dividends or stock splits, providing a more accurate long-term price history.
volume CoinAPI [2023] indicating the level of trading activity and movement, typically the change in price from the previous day, reflects short-term price fluctuations.
devived financial indicators
2 movement Tardi [2022], Jegadeesh [1990], Jegadeesh and Titman [1993] represents the change in stock price between consecutive trading days.
returns Tardi [2022], Jegadeesh [1990], Jegadeesh and Titman [1993] represents the percentage change in the closing price over consecutive days.
VWAP (Volume Weighted Average Price) Mitchell et al. [2020] calculates the average price of a stock weighted by trading volume, providing a benchmark for institutional investors.
3 macd_histogram (values) Anghel [2015] measures the difference between a stock’s MACD line and its signal line, helping identify trends and momentum.
macd_crossover (signals) Anghel [2015] indicates a buy or sell signal when the MACD line crosses above or below the signal line.
4 bollinger_bands (signals) Bollinger [2001] indicates whether the adjusted close price exceeds the upper or lower price limits around a moving average, serving as a measure of price volatility.
exceeding_upper/lower (values) Bollinger [2001] the value of the adjusted close price exceeds the upper or lower price limits around a moving average.
5 overbought_and_oversold_conditions Wilder [1978] classify the market state into "overbought" or "oversold" areas based on predefined thresholds for $K$, $D$, and $J$.
kdj_crossover (signals) Wu and Diao [2015] identify "bullish" or "bearish" signals based on the relationship between $K$ and $D$.
6 alpha_smr Kakushadze [2016] measures short-term reversal patterns by capturing stocks’ tendency to revert after abrupt price movements.
alpha_mom Kakushadze [2016] tracks momentum effects by identifying stocks with strong recent performance likely to continue trending.
alpha_009 Kakushadze [2016] uses recent price changes to classify trends based on their range within a 5-day window.
alpha_012 Kakushadze [2016] relates volume change direction to the negative change in price.
alpha_041 Kakushadze [2016] measures the difference between price range midpoint and VWAP.
alpha_049 Kakushadze [2016] detects strong downtrends or reversals in lagged price movements.
alpha_051 Kakushadze [2016] quantifies mean-reversion behavior over medium horizons, typically exploiting price corrections after sustained moves.
alpha_054 Kakushadze [2016] combines momentum and volume signals to detect stocks with improving trends supported by trading activity.
alpha_101 Kakushadze [2016] normalizes price change by daily price range.
7 alpha_002 Kakushadze [2016] measures the negative 6-day correlation between ranked changes in logarithmic volume and ranked price movements.
alpha_006 Kakushadze [2016] computes the negative correlation between opening price and volume over a 10-day window.
alpha_024 Kakushadze [2016] compares long-term and short-term price movements to identify breakouts or retracements.
alpha_028 Kakushadze [2016] scales the relationship between average price, low price correlation, and close price.
alpha_032 Kakushadze [2016] combines deviations from a 7-day moving average and long-term correlation between VWAP and lagged close.
alpha_053 Kakushadze [2016] identifies stocks exhibiting strong recent performance (typically over 3-6 months) while controlling for volatility, aiming to capture persistent trends while mitigating risk from erratic price swings.
8 alpha_021 Kakushadze [2016] identifies trends based on short- and long-term price averages and volume conditions.
alpha_023 Kakushadze [2016] tracks changes in high prices if they exceed their 20-day average.
alpha_046 Kakushadze [2016] analyzes trends in lagged and current 10-day price changes.
E.3 Sequences and Serialization
An example in our datastore is shown below.
"stock_name":"ABBV",
"query_date":"2014-06-05",
"movement":"rise",
"open":54.549999,
"high":55.32,
"low":54.360001,
"close":55.299999,
"adj_close":36.980961,
"volume":4847300,
"macd_histogram":0.0621141602,
"macd_crossover":null,
"bollinger_bands":null,
"exceeding_upper":null,
"exceeding_lower":null,
"overbought_and_oversold_conditions":null,
"kdj_crossover":null,
"returns":0.0131913094,
"VWAP":54.9933333333,
"alpha_smr":0.000549858,
"alpha_mom":0.0121661224,
"alpha_002":null,
"alpha_006":null,
"alpha_009":-0.481476,
"alpha_012":-0.481476,
"alpha_021":1,
"alpha_023":0.0,
"alpha_024":-0.769035,
"alpha_028":null,
"alpha_032":null,
"alpha_041":-0.1554335256,
"alpha_046":-0.481476,
"alpha_049":-0.481476,
"alpha_051":-0.481476,
"alpha_053":null,
"alpha_054":126.4264005641,
"alpha_101":-18.282056485
Appendix F Experimental Settings
F.1 StockLLM-1B-Instruct Backbone
To activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA 3.2-1B-Instruct) [Grattafiori et al., 2024] using the LoRA technique for efficient low-rank adaptation, resulting in StockLLM-1B-Instruct. By intentionally using a smaller backbone model, we establish a more challenging experimental setup, ensuring that performance improvements are attributable to FinSeer’s retrieval capabilities rather than the LLM’s capacity. More details of fine-tuning are illustrated in Appendix F.1.
The fine-tuning process is implemented using the LlamaFactory framework [Zheng et al., 2024], with the following configuration: a learning rate of 5e-5, a cosine scheduler, gradient accumulation over 8 steps, mixed-precision (fp16) training, and 5 epochs of training with regular evaluation to log metrics and select the best-performing checkpoint. This setup demonstrates the robustness of our framework in extracting meaningful insights even under constrained model size, highlighting the effectiveness of FinSeer in enhancing financial forecasting tasks. The fine-tuning prompt is shown in Figure 5, with candidate sequences randomly retrieved from the dataset. The fine-tuning step is to train the StockLLM-1B-Instruct to follow instructions, without additional steps for retrieving certain candidates.
F.2 Experimental Settings
Baselines. The introduction of retrieval baselines is as follows:
  • Random retrieval: The candidate range is aligned with the "Candidate Pool Settings" part. We use the random.sample() function in Python to retrieve random candidates from the candidate pool.
  • DTW (Dynamic Time Warping) distance: The formula measures the similarity between two temporal sequences by aligning them non-linearly in time, minimizing the cumulative warping cost. $D_{\text{DTW}}(X,Y)=\min_{\pi\in\mathcal{P}}\sqrt{\sum_{(i,j)\in\pi}(x_{i}-y_{j})^{2}}$, where $X,Y$ are sequences, $\pi$ is a warping path, and $\mathcal{P}$ is the set of all possible paths.
  • Instructor-large [Su et al., 2023]: A 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors.
  • UAE-large-v1 [Li and Li, 2023]: A 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors.
  • E5-mistral-7b-instruct [Wang et al., 2023]: A 7111M embedder initialized from Mistral-7B-v0.1 [Jiang et al., 2023] and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
  • BGE-large-en-v1.5 [Xiao et al., 2023]: A 335M general embedder pre-trained with RetroMAE [Xiao et al., 2022], encoding sequences into 1024-dimensional tensors.
  • LLM-Embedder [Zhang et al., 2023]: A 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs. It encodes sequences into 768-dimensional tensors.
Prompts. The prompt for bare StockLLM-1B-Instruct is shown in Figure 6, and the prompt for retrieval methods is shown in Figure 5.
Figure 5: The prompt template for StockLLM-1B-Instruct fine-tuning and RAG methods evaluation.
Figure 6: The prompt template for bare StockLLM-1B-Instruct evaluation.
Hyperparameters. For each model, we employ default parameter settings, utilizing official models for open-source LLMs obtained from Hugging Face. These testing procedures take place on a computational infrastructure consisting of one NVIDIA A800 Tensor Core GPU, equipped with 80GB of memory.
Error Bars. To account for potential variability, we conducted three randomized retrieval trials and report averaged performance metrics. For similarity-based retrieval methods, since both queries and candidates are encoded as fixed vector representations, the retrieved candidates remain deterministic across runs. To verify this stability, we performed three identical inference passes through StockLLM, observing minimal variance in prediction outcomes.
Appendix G Addtional Experimental Results
G.1 Candidate Movement Correlation
@@
In this section, we investigate whether StockLLM-1B-Instruct relies on the movement trends of retrieved sequences to make predictions. To this end, we compute the correlation between the movements of retrieved sequences and the LLM’s generated results.
@@
The results in Table 5 reveal no significant correlation between the movement direction (rise or fall) of the retrieved candidate sequences and the final predictions of StockLLM. This finding indicates that StockLLM-1B-Instruct does not simply mirror the movement trends of the retrieved sequences but instead analyzes their specific content to infer the query’s movement direction. This ability highlights the LLM’s capacity to extract meaningful insights from complex time-series data.
@@
Table 5: Correlation between the movements of retrieved sequences and LLM’s generated results.
@@
This observation underscores the critical role of retrieval quality in RAG models, as the relevance and informativeness of retrieved sequences directly influence prediction performance. Notably, StockLLM-1B-Instruct is a relatively small 1B parameter LLM, making this a particularly challenging setting. Despite this, our time-series-tailored RAG model, FinSeer, successfully enhances StockLLM’s performance by retrieving rich and relevant time-sensitive information. This demonstrates the effectiveness of our approach in improving prediction accuracy and showcases its potential for financial time-series forecasting.
G.2 Indicator Sequence Visualization
@@
To intuitively examine how FinSeer embeds indicator sequences, we use LargeVis [Tang et al., 2016] to reduce the embeddings to two dimensions. In Figure 7, we visualize [^6] https://medviz.org/app/ the vector space of nine indicators, including adjusted close price, open price, high price, low price, close price, volume, movement, returns, and WAP. The results show that the green points, representing movement information, are well-clustered, indicating that FinSeer effectively captures meaningful patterns. However, this also suggests that deeper relationships between stock movements and indicators remain to be explored, highlighting the potential for further research in this area.
Figure 7: Visualization of indicator sequence embeddings on ACL18 dataset.
G.3 Results of StockLLM-8B-Instruct
We additionally trained StockLLM-8B-Instruct based on LLaMA-3.1-8B-Instruct to evaluate FinSeer’s performance on larger LLMs. As shown in Table 6, the results demonstrate that FinSeer can consistently enhance LLM performance regardless of the backbone model’s inherent capability for stock movement prediction. Compared to the 1B version, we observe greater performance variability in the bare LLM, while different retrieval methods maintain similar relative impacts on StockLLM’s performance. Notably, random retrieval and other retrieval methods continue to show inconsistent effects, sometimes improving but other times degrading performance.. FinSeer not only outperforms all retrieval-based approaches but also delivers stable performance enhancements to StockLLM, validating the effectiveness of our RAG framework.
Table 6: Results of stock movement predictions using StockLLM-8B-Instruct and retrieval models.
Retrieving Methods ACL18 BIGDATA22 STOCK23
(+ StockLLM-8B-Instruct) ACC MCC ACC MCC ACC MCC
w/o Retrieval 0.479 -0.041 0.446 -0.108 0.510 0.020
Random Retrieval 0.509 0.028 0.450 -0.149 0.495 -0.012
DTW 0.509 0.020 0.457 -0.087 0.496 -0.008
Instructor 0.497 -0.005 0.457 -0.087 0.506 0.011
UAE 0.509 0.019 0.465 -0.081 0.509 0.018
E5 0.484 -0.031 0.453 -0.099 0.498 -0.005
BGE 0.522 0.058 0.460 -0.104 0.507 -0.001
LLM Embedder 0.486 -0.034 0.454 -0.123 0.499 -0.002
FinSeer 0.554 0.122 0.469 -0.065 0.511 0.023